OpenStudy (anonymous):

A tea bag is shaped like a regular square pyramid. Each leg of the base is 4 cm, and the slant height is 5 cm. What are the lateral area & surface area of the tea bag?

OpenStudy (anonymous):

40cm^3

OpenStudy (anonymous):

LA of pyramid = (1/2)(perimeter of base)(slant height) LA=(1/2)(16cm)(5cm)=40cm^2 Total Surface area of pyramid = (1/2)(perimeter of base)(slant height)+(area of base) TSA=(1/2)(16cm)(5cm)+(4 cm)^2=44 cm^2